# SAMSGL: Series-Aligned Multi-Scale Graph Learning for Spatio-Temporal Forecasting

### v2

@@ Abstract
Index Terms:
Spatio-temporal forecasting, time series forecasting, graph nearual network, graph structure learning.
@@ I Introduction
Spatio-temporal forecasting plays a pivotal role in predicting forthcoming states within spatio-temporal networks across diverse domains like traffic prediction [65, 66, 67], weather forecasting [2, 9, 64] and power grid prediction [56, 75]. Despite these achievements, the performance of spatio-temporal forecasting hinges on two crucial factors: 1) time delays in propagation dynamics: it is time-consuming to exchange messages between spatial nodes and the propagation speed during spatial interactions acts as a determinant affecting inter-node time delays. It is contingent not only upon spatial distance, but also intricacies like conditions and directions of propagation paths among nodes. For instance, in traffic forecasting, holidays impact road conditions, influencing traffic speed [48] and leading to increased time delays between nodes. Similarly, shifts in global weather patterns, such as monsoons in meteorological forecasting [1], alter propagation direction and speed, affecting time delays. 2) high-dimensional interactions: the complexity of high-dimensional networks characterized by a large number of nodes and the complex connections between them [74] presents a formidable challenge in capturing inter-node interactions, thereby intensifying the intricacy of spatio-temporal forecasting [3]. In particular, traffic flow forecasting employs thousands of sensors [45], and high-resolution weather forecasting involves tens of thousands of grids [7]. This intricate interplay extends beyond proximate neighbors, incorporating associations between distant and adjoining nodes, thus complicating the capture of spatio-temporal interactions among nodes.
@@
While approaches like Graph Convolution Networks (GCN) [36] endeavor to simulate node propagation dynamics by aggregating spatial information from neighboring nodes, time delays among nodes introduce a challenge in accurately capturing propagation dynamics. However, GCN-based methods aggregate messages from nodes within the same time step, without considering the influence of node features from previous time steps that genuinely impact the present features of nodes. To address the influence of interaction time delays and better learn propagation dynamics, some series alignment methods have emerged. For instance, Corrformer [9] aligns node series prior to computing cross-correlation using a tree structure. Although Corrformer embeds node locations, it does not fully account for spatial interaction factors, such as distance, propagation conditions, and propagation path direction, which affect time delays. Consequently, there is a pressing need to develop an approach that effectively aggregates messages from nodes’ spatially relevant features, while reducing the influence of time delays.
@@
Graph structure learning methods have been developed to depict the high-dimensional intricate interactions between nodes. The structure of a graph is learned based on nodes’ similarity [27] or generated with trainable node embeddings. For example, Regularized Graph Structure Learning (RGSL) [30] captures node interactions from implicit and prior graphs. Given the complexity of high-dimensional interactions, the representation of implicit interactions with a single graph may be insufficient to successfully address this issue. Therefore, employing multi-scale graphs that portray interactions at both local explicit and global implicit scales holds significant promise to tackle these intricate high-dimensional interactions [17].
@@
Furthermore, to capture spatio-temporal interactions among nodes, aside from the construction of multi-scale graph structures, it is essential to employ efficient spatio-temporal architectures. While there have been notable advancements in this regard, such as the development of Graph Convolution Recurrent Neural Networks (GCRNN) [13, 16] and the integration of graph convolution with attention mechanisms [18, 47], there remains potential for developing a spatio-temporal architecture that can fuse spatial and temporal information more effectively and achieve better forecasting performance.
@@
  • Numerous experiments on real-world datasets for meteorology forecasting and traffic flow forecasting tasks are conducted to demonstrate the superior performance compared to existing models, as shown in Fig. 2 and other parts.
@@ II Related Work
Graph convolution which aims to aggregate features and propagate messages from nodes, can be categorized into two types: spectral and spatial graph convolution. Spectral graph convolution can be conceptualized as filtering the graph signals, where the filtering operation of graph can be realized as a self-loop Laplacian matrix [32] or approximated by Chebyshev polynomials [33]. Efforts to enhance graph representation include refining the graph transformation process for graph signals. In [11], AKGNN is introduced as a data-driven graph kernel learning module, dynamically adjusting the trade-off between all-pass and low-pass filters. To address instability in feature transformation, [34] proposes Ortho-GConv, ensuring robust orthogonality. Additionally, [35] introduces a network enforcing invariance in the sign and basis of eigenvectors, thereby enhancing the expression of spectral graph convolution.
@@
Spatial graph convolution aims to replicate the message propagation process [72] and aggregates information from neighboring nodes [73]. [51] proposes an aggregation strategy that combines features from lower to higher-order neighborhoods in a non-recursive way by employing a randomized path exploration approach. For tasks unsuitable for Euclidean space, some methods map the graph to alternative spaces. [39] utilizes Riemannian manifolds to represent spatial correlations between stocks. SelfMGNN [38] constructs a mixed-curvature space via the Cartesian product of multiple Riemannian component spaces to learn and fuse graph representations across these spaces. These methods primarily focus on information aggregation at a single time step, with less emphasis on considering time delays in message propagation. Our series-aligned graph convolution also aims to achieve better aggregation of node features by performing convolutions on aligned time series, thereby reducing the influence of time delays.
II-B Graph Structure Learning.
@@
In spatio-temporal forecasting, graph structure learning is employed to establish suitable relationships between nodes, thereby enhancing spatio-temporal representation. Within graph structure learning, some methods directly create a graph and employ it in downstream tasks, while others concentrate on refining existing structures derived from similarities or physical connections. Some approaches directly generate an adjacency matrix from node sequences with node similarity. For instance, STFGNN [27] employs dynamic time warping to measure node sequence similarity. ST-LGSL [28] utilizes latent graph topological information based on MLP-KNN. To tackle graph structure learning of large scale graph or long sequences, some researches use neural networks to learn a graph. Other approaches involve generating graphs from node embeddings. For instance, RGSL [30] captures implicit graph information by a regularized graph generation module. Additionally, the Meta-Graph Convolutional Recurrent Network (MGCRN) [16] uses hyper-network and node embedding banks to disentangle graphs in both space and time. The Possibilistic Neighborhood Graph (PNG) [52] adaptively measures the possibility of two nodes being neighbors.
@@
To refine the original graph structure, AdaSTN [29] employs temporal convolution to assess sequence similarity, using the learned graph to fine-tune the original structure. [61] optimizes the graph iteratively by refining the structure with graphs generated from similarity metrics. Establishing better relations between nodes, graph structure learning has demonstrated its effectiveness in enhancing graph representation. Therefore, in this paper, we focus on constructing multi-scale graph structure learning that generates graphs from global delayed and non-delayed scales, as well as local scales to capture remote and neighboring node interactions.
II-C Spatio-temporal Architecture.
@@
Spatio-temporal forecasting tasks can be approached with or without considering spatial information. When treating all node states as variables and disregarding location information, multivariate forecasting techniques [69] are employed to to predict future dynamics. These methods encompass models like the autoregressive integrated moving average model [21], Recurrent Neural Networks (RNN) [22], and transformer-based structures [9]. They tend to focus on capturing the seasonal and trend information within time series [56] or decompose time series into various frequency components [70]. Some approaches emphasize temporal evolution and variable characteristics, aiming to capture relations between variables [57] or accommodate variable-specific shifts [59]. In contrast, efforts have also been made to incorporate location information into multivariate methods. Techniques like embedding spatial positional formation into features have been explored. For instance, [9] embeds the loaction information into the features. By spectral transformation of the graph structure, [60] embeds the eigenvalues into features to make the network aware of the structure. Additionally, graph-based methods are utilized for representing spatial relationships.
@@
Furthermore, graph-based methods employ adjacency matrices to illustrate the connections between nodes, encompassing factors such as road connectivity and distance between nodes. Some approaches combine graphs with RNN by incorporating graph convolution features into both input and hidden states. For instance, TGCN [25] combines combined GCN and the gated recurrent unit to capture the spatio-temporal dependency. Different type of graph convolution layers are applied for better spatial representation. DCRNN [24] applied a diffusion convolution layer model the traffic diffusion process, while AGCRN [13] proposed a node and data adaptive graph convolution to achieve the node specific graph convolution. Others benefit from alternating graph networks and temporal modules, as well as stacked multi-layer spatio-temporal blocks. ASTGCN [18] use a spatial temporal attention to predict traffic flow by alternately learning the spatial and temporal attention modules. Approaches that consider both spatial and temporal relationships typically yield improved performance in spatio-temporal forecasting tasks, emphasizing the significance of effectively combining spatial and temporal interactions. In this paper, we propose a spatio-temporal architecture that combines multi-scale graph structure learning and spatio-temporal fusion Graph-FC blocks to learn the complex spatio-temporal dependencies.
@@ III Methodologies
In this section, we present the framework of the proposed SAMSGL, as illustrated in Fig. 1. We begin by defining the problem and introducing the notations in this study. Subsequently, we describe the key components of our model,which includes: the series-aligned Graph Convolution which capture the time delay node features, the multi-scale graph structure learning module that obtains global delayed and non-delayed features and local features, and Graph-FC blocks that combine the spatial and temporal information. The overall workflow of our SAMSGL is depicted in Fig. 1(a), encompassing the multi-scale graph structure learning and the Graph-FC blocks. The specifics of multi-graph convolution in Graph-FC blocks are outlined in Fig. 1(b), while the series-aligned graph convolution workflow is presented in Fig. 1(c).
Figure 1: The framework of Series-Aligned Multi-Scale Graph Learning (SAMSGL). (a) The overall structure of SAMSGL. (b) The structure of Multi-Graph Convolution. (c) The workflow of Series-Aligned Graph Convolution where fft represents fast Fourier transform and ifft is inversed fast Fourier transform.
III-A Preliminaries
Spatio-temporal forecasting involves predicting future dynamics of spatio-temporal systems. In graph-based models, the interactions between $N$ nodes are represented by the adjacent matrix $\mathcal{A}\in\mathbb{R}^{{N}\times{N}}$, where the elements of the adjacent matrix indicate interaction coefficients among nodes. The state of node $i$ at time step $t$ is denoted as $X_{i,t}\in\mathbb{R}^{C}$, and $C$ is the state’s dimensions for each node. For spatio-temporal forecasting tasks, the input comprises historical time series with length $L$ of $N$ nodes, which can be represented as $X\in\mathbb{R}^{{N}\times{L}\times{C}}$. The objective is to train a network $\mathcal{F}_{\theta}$ with parameters $\theta$ to predict future series $X^{{}^{\prime}}\in\mathbb{R}^{{N}\times{L^{{}^{\prime}}}\times{C}}$ with length $L^{{}^{\prime}}$ as follows:
@@ III-A Preliminaries
III-B Series-Aligned Graph Convolution
Due to the existence of time delays in propagation dynamics, the states of target nodes may be truly influenced by the states of its neighboring nodes a few time steps ago. However, traditional graph convolution aggregates state messages of neighboring nodes at the same time step, without considering time delays in propagation. To illustrate, due to time delays in propagation, the state $X_{i,t}$ of node $i$ at time step $t$ may be influenced not only by its immediate previous neighboring node’s state $X_{i-1,t-1}$, but also by the state of its neighboring node $X_{j,t-\tau_{i,j}}$, where $\tau_{i,j}$ represents the delay between node $i$ and node $j$. Therefore, it is necessary to consider time delays by series alignment [9] before graph convolution. Our series-aligned graph convolution aims to enhance the aggregation of related node features by conducting series alignment, thereby mitigating the impact of intricate time delays between nodes.
We firstly implement a series alignment module to ascertain time delays between nodes. Recognizing that calculating the delay between each node pair is computationally demanding, we set a reference node $\bar{q}$ and calculate the delay of each node relative to the reference node. By aligning all the nodes with the reference node, we achieve computational efficiency while still obtaining relevant delay information. To accomplish this, inspired by [9], we identify the time step with the highest correlation as the initial point of the time series of each node and align the series. The process unfolds as follows:

### v3

@@ Abstract
preprint: AIP/123-QED
Accurate prediction of spatio-temporal systems is crucial for understanding system conditions and making effective decisions. This paper introduces the Series-Aligned Multi-Scale Graph Learning (SAMSGL) framework, aimed at capturing propagation dynamics and modeling high-dimensional interactions. Firstly, we propose a series-aligned graph convolution layer to address time delays between nodes, thereby enhancing spatial message aggregation. Additionally, to extract spatio-temporal interactions from data, we present a comprehensive spatio-temporal architecture incorporating a multi-scale graph structure learning module and Graph-fully connected (Graph-FC) blocks. Within the multi-scale graph structure learning module, adaptive node embeddings and node locations are utilized to derive global spatial delayed and non-delayed relations, as well as local information. The Graph-FC blocks effectively fuse spatial information from multiple graphs and temporal information. Experimental results demonstrate that our model improves prediction accuracy compared to previous approaches. We validate the effectiveness of our modules through ablation studies.
@@ I Introduction
Spatio-temporal forecasting plays a pivotal role in predicting forthcoming states within spatio-temporal networks across diverse domains like chaotic prediction Liu et al. (2024); Wang et al. (2024); Aksoy (2024), traffic prediction Kwon, Jung, and Eom (2023), weather forecasting Wu et al. (2023); Cheng, Guo, and Arcucci (2023) and power grid prediction Titz et al. (2024); Xiong et al. (2022). Despite these achievements, the performance of spatio-temporal forecasting hinges on two crucial factors: 1) time delays in propagation dynamics: it is time-consuming to exchange messages between spatial nodes and the propagation speed during spatial interactions acts as a determinant affecting inter-node time delays. It is contingent not only upon spatial distance, but also intricacies like conditions and directions of propagation paths among nodes. For instance, in traffic forecasting, holidays impact road conditions, influencing traffic speed Macioszek and Kurek (2021) and leading to increased time delays between nodes. Similarly, shifts in global weather patterns, such as monsoons in meteorological forecasting Boers et al. (2019), alter propagation direction and speed, affecting time delays. 2) high-dimensional interactions: the complexity of high-dimensional networks characterized by a large number of nodes and the complex connections between them Tang et al. (2020) presents a formidable challenge in capturing inter-node interactions, thereby intensifying the intricacy of spatio-temporal forecasting Ji et al. (2023). In particular, traffic flow forecasting employs thousands of sensors Chen et al. (2001), and high-resolution weather forecasting involves tens of thousands of grids Lam et al. (2023). This intricate interplay extends beyond proximate neighbors, incorporating associations between distant and adjoining nodes, thus complicating the capture of spatio-temporal interactions among nodes. Integrating time delays in propagation dynamics and capturing high-dimensional interactions tend to enhance prediction accuracy.
@@
While approaches like Graph Convolution Networks (GCN) Kipf and Welling (2017) endeavor to simulate node propagation dynamics by aggregating spatial information from neighboring nodes, time delays among nodes introduce a challenge in accurately capturing propagation dynamics. However, GCN-based methods aggregate messages from nodes within the same time step, without considering the influence of node features from previous time steps that genuinely impact the present features of nodes. To address the influence of interaction time delays and better learn propagation dynamics, some series alignment methods have emerged. For instance, Corrformer Wu et al. (2023) aligns node series prior to computing cross-correlation using a tree structure. Although Corrformer embeds node locations, it does not fully account for spatial interaction factors, such as distance, propagation conditions, and propagation path direction, which affect time delays. Consequently, there is a pressing need to develop an approach that effectively aggregates messages from nodes’ spatially relevant features, while reducing the influence of time delays.
@@
Graph structure learning methods have been developed to depict the high-dimensional intricate interactions between nodes. The structure of a graph is learned based on nodes’ similarity Li and Zhu (2021) or generated with trainable node embeddings. For example, Regularized Graph Structure Learning (RGSL) Yu et al. (2022) captures node interactions from implicit and prior graphs. Given the complexity of high-dimensional interactions, the representation of implicit interactions with a single graph may be insufficient to successfully address this issue. Therefore, employing multi-scale graphs that portray interactions at both local explicit and global implicit scales holds significant promise to tackle these intricate high-dimensional interactions Shao et al. (2022).
@@
Furthermore, to capture spatio-temporal interactions among nodes, aside from the construction of multi-scale graph structures, it is essential to employ efficient spatio-temporal architectures. While there have been notable advancements in this regard, such as the development of Graph Convolution Recurrent Neural Networks (GCRNN) Bai et al. (2020); Jiang et al. (2023a) and the integration of graph convolution with attention mechanisms Guo et al. (2019); Li et al. (2022), there remains potential for developing a spatio-temporal architecture that can fuse spatial and temporal information more effectively and achieve better forecasting performance.
@@
  • Numerous experiments on real-world datasets for meteorology forecasting and traffic flow forecasting tasks are conducted to demonstrate the superior performance compared to existing models.
@@ II Related Work
Graph convolution which aims to aggregate features and propagate messages from nodes, can be categorized into two types: spectralBruna et al. (2014); Defferrard, Bresson, and Vandergheynst (2016); Ju et al. (2022); Guo et al. (2022) and spatial graph convolutionJiang et al. (2023b); Zhang, Wu, and Yan (2023); Bose and Das (2023); Sawhney et al. (2021); Guo et al. (2024). Spectral graph convolution can be conceptualized as filtering the graph signals, where the filtering operation of graph can be realized as a self-loop Laplacian matrix Bruna et al. (2014) or approximated by Chebyshev polynomials Defferrard, Bresson, and Vandergheynst (2016). Efforts to enhance graph representation include refining the graph transformation process for graph signals. AKGNNJu et al. (2022) is introduced as a data-driven graph kernel learning module, dynamically adjusting the trade-off between all-pass and low-pass filters. To address instability in feature transformation, Guo et al.Guo et al. (2022) propose Ortho-GConv, ensuring robust orthogonality.
@@
Spatial graph convolution aims to replicate the message propagation process Jiang et al. (2023b) and aggregates information from neighboring nodes Zhang, Wu, and Yan (2023). Bose et al.Bose and Das (2023) propose an aggregation strategy that combines features from lower to higher-order neighborhoods in a non-recursive way by employing a randomized path exploration approach. For tasks unsuitable for Euclidean space, some methods map the graph to alternative spaces. Sawhney et al. Sawhney et al. (2021) utilize Riemannian manifolds to represent spatial correlations between stocks. These methods primarily focus on information aggregation at a single time step, with less emphasis on considering time delays in message propagation. Our series-aligned graph convolution also aims to achieve better aggregation of node features by performing convolutions on aligned time series, thereby reducing the influence of time delays.
II.2 Graph Structure Learning.
@@
In spatio-temporal forecasting, graph structure learning is employed to establish suitable relationships between nodes, thereby enhancing spatio-temporal representation. Within graph structure learning, some methods directly create a graph and employ it in downstream tasks, while others concentrate on refining existing structures derived from similarities or physical connections. Some approaches directly generate an adjacency matrix from node sequences with node similarity. For instance, STFGNN Li and Zhu (2021) employs dynamic time warping to measure node sequence similarity. To tackle graph structure learning of large scale graph or long sequences, some researches use neural networks to learn a graph. Other approaches involve generating graphs from node embeddings. For instance, RGSL Yu et al. (2022) captures implicit graph information by a regularized graph generation module. Additionally, the Meta-Graph Convolutional Recurrent Network (MGCRN) Jiang et al. (2023a) uses hyper-network and node embedding banks to disentangle graphs in both space and time.
@@
To refine the original graph structure, AdaSTN Ta et al. (2022) employs temporal convolution to assess sequence similarity, using the learned graph to fine-tune the original structure. Chen et al.Chen, Wu, and Zaki (2020) optimize the graph iteratively by refining the structure with graphs generated from similarity metrics. Establishing better relations between nodes, graph structure learning has demonstrated its effectiveness in enhancing graph representation. Therefore, in this paper, we focus on constructing multi-scale graph structure learning that generates graphs from global delayed and non-delayed scales, as well as local scales to capture remote and neighboring node interactions.
II.3 Spatio-temporal Architecture.
@@
Spatio-temporal forecasting tasks can be approached with or without considering spatial information. When treating all node states as variables and disregarding location information, multivariate forecasting techniques Zhang et al. (2023); Feng et al. (2024) are employed to to predict future dynamics. These methods encompass models like the autoregressive integrated moving average model Kothapalli and Totad (2017), Recurrent Neural Networks (RNN) Shao and Soong (2016), and transformer-based structures Wu et al. (2023). They tend to focus on capturing the seasonal and trend information within time series Rebei, Amayri, and Bouguila (2023); Jiang et al. (2024) or decompose time series into various frequency components Wu et al. (2022). Some approaches emphasize temporal evolution and variable characteristics, aiming to capture relations between variables Zhang and Yan (2022) or accommodate variable-specific shifts Cirstea et al. (2022). In contrast, efforts have also been made to incorporate location information into multivariate methods. Techniques like embedding spatial positional formation into features have been explored. For instance, by spectral transformation of the graph structure, Jiang et al. (2023c) embeds the eigenvalues into features to make the network aware of the structure. Additionally, graph-based methods are utilized for representing spatial relationships.
@@
Furthermore, graph-based methods Zou et al. (2024); Nauck et al. (2023); Zhao et al. (2020); Li et al. (2018); Bai et al. (2020) employ adjacency matrices to illustrate the connections between nodes, encompassing factors such as road connectivity and distance between nodes. Some approaches combine graphs with RNN by incorporating graph convolution features into both input and hidden states. For instance, TGCN Zhao et al. (2020) combines combined GCN and the gated recurrent unit to capture the spatio-temporal dependency. Different type of graph convolution layers are applied for better spatial representation. DCRNN Li et al. (2018) applied a diffusion convolution layer model the traffic diffusion process, while AGCRN Bai et al. (2020) proposed a node and data adaptive graph convolution to achieve the node specific graph convolution. Others benefit from alternating graph networks and temporal modules, as well as stacked multi-layer spatio-temporal blocks. ASTGCN Guo et al. (2019) use a spatial temporal attention to predict traffic flow by alternately learning the spatial and temporal attention modules. Approaches that consider both spatial and temporal relationships typically yield improved performance in spatio-temporal forecasting tasks, emphasizing the significance of effectively combining spatial and temporal interactions. In this paper, we propose a spatio-temporal architecture that combines multi-scale graph structure learning and spatio-temporal fusion Graph-FC blocks to learn the complex spatio-temporal dependencies.
@@ III Methodologies
In this section, we present the framework of the proposed SAMSGL, as illustrated in Fig. 1. We begin by defining the problem and introducing the notations in this study. Subsequently, we describe the key components of our model,which includes: the series-aligned Graph Convolution which captures the time delay node features, the multi-scale graph structure learning module that obtains global delayed and non-delayed features and local features, and Graph-FC blocks that combine the spatial and temporal information. The overall workflow of our SAMSGL is depicted in Fig. 1(a), encompassing the multi-scale graph structure learning and the Graph-FC blocks. The specifics of multi-graph convolution in Graph-FC blocks are outlined in Fig. 1(b), while the series-aligned graph convolution workflow is presented in Fig. 1(c).
Figure 1: The framework of Series-Aligned Multi-Scale Graph Learning (SAMSGL). (a) The overall structure of SAMSGL. The Multi-scale graph structure learning module is developed to generate spatial relations of different scales, which are denoted as $\mathcal{A}_{\rm nd}$, $\mathcal{A}_{\rm d}$, and $\mathcal{A}_{\rm l}$, respectively. These generated adjacency matrices, along with the node states $X$, are input into stacks of Graph-FC blocks. (b) The structure of Multi-Graph Convolution in the Graph-FC blocks. (c) The workflow of Series-Aligned Graph Convolution where "fft" represents fast Fourier transform and "ifft" is inversed fast Fourier transform.
III.1 Preliminaries
Spatio-temporal forecasting involves predicting future dynamics of spatio-temporal systems. In graph-based models, the interactions between $N$ nodes are represented by the adjacency matrix $\mathcal{A}\in\mathbb{R}^{{N}\times{N}}$, where the elements of the adjacency matrix indicate interaction coefficients among nodes. The state of node $i$ at time step $t$ is denoted as $X_{i,t}\in\mathbb{R}^{C}$, and $C$ is the state’s dimensions for each node. For spatio-temporal forecasting tasks, the input comprises historical time series with length $L$ of $N$ nodes, which can be represented as $X\in\mathbb{R}^{{N}\times{L}\times{C}}$. The objective is to train a network $\mathcal{F}_{\theta}$ with parameters $\theta$ to predict future series $X^{{}^{\prime}}\in\mathbb{R}^{{N}\times{L^{{}^{\prime}}}\times{C}}$ with length $L^{{}^{\prime}}$ as follows:
@@ III.1 Preliminaries
III.2 Series-Aligned Graph Convolution
Due to the existence of time delays in propagation dynamics, the states of target nodes may be truly influenced by the states of its neighboring nodes a few time steps ago. However, traditional graph convolution aggregates state messages of neighboring nodes at the same time step, without considering time delays in propagation. To illustrate, due to time delays in propagation, the state $X_{i,t}$ of node $i$ at time step $t$ may be influenced not only by its immediate previous neighboring node’s state $X_{i-1,t-1}$, but also by the state of its neighboring node $X_{j,t-\tau_{i,j}}$, where $\tau_{i,j}$ represents the delay between node $i$ and node $j$. Therefore, it is necessary to consider time delays by series alignment Wu et al. (2023) before graph convolution. Our series-aligned graph convolution aims to enhance the aggregation of related node features by conducting series alignment, thereby mitigating the impact of intricate time delays between nodes. The procedure, as shown in Fig. 1 (c), involves a sequence of steps: series alignment, followed by graph convolution, and finally, an aligning back operation.
We firstly implement a series alignment module to ascertain time delays between nodes. Recognizing that calculating the delay between each node pair is computationally demanding, we set a reference node $\bar{q}$ and calculate the delay of each node relative to the reference node. By aligning all the nodes with the reference node, we achieve computational efficiency while still obtaining relevant delay information. To accomplish this, inspired by Wu et al.Wu et al. (2023), we identify the time step with the highest correlation as the initial point of the time series of each node and align the series. The process unfolds as follows: